OpenStudy (anonymous):

Give the slope and y-intercept for the equation below. 3x - 4y = -2 m =? y-intercept =?

OpenStudy (aroub):

-4y=-3x-2 y=3/4x+1/2 slope=3/4 y-intercept=1/2

OpenStudy (anonymous):

get it in the form of y = mx+b -4y = -3x-2 Multiply everything by (-1) 4y= 3x+2 Divide by 4 y =3/4x =1/2 m= 3/4 y-int= 1/2